OpenStudy (anonymous):

whats 2x^2+7x-18 A.(x-3)(x-6) b.(x-3)(x+6) c.(x+2)(x-9) D.(x-2)(x+9)

OpenStudy (anonymous):

try multiplying the possible answers and see which one gives you 2x^2+7x-18

OpenStudy (anonymous):

tha answer is A!

OpenStudy (amistre64):

the clues given by the equation are: 2x^2(+)7x (-)18 ^ ^ ^ +*- = - ; so the signs are different ^ larger factor of (18) is positive

OpenStudy (amistre64):

+6-3 = NOT7 so thats out +9-2 = 7, so its gonna be one of the options that fits the bill

OpenStudy (amistre64):

determine which option has a +9 and a -2 in it, and go with it :)

OpenStudy (anonymous):

can u help me